OpenStudy (anonymous):

Find the supplement of the complement of an angle measuring 29°.

OpenStudy (anonymous):

supplement = 180 - 29 complement = 90 - 29

OpenStudy (anonymous):

Ive tried findng the answer by doing it those two ways....the site im doing my homework on says its incorrect though